OpenStudy (anonymous):

. Imagine you are teaching a fellow student how to solve: 2(x + 6) − 10 = 12 + 4(x − 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

First you distribute: 2x+12-10=12+4x-4 Then you combine like terms: 2x+2=4x+8 Next you subtract the x's from one side: 2=2x+8 Subract the number alone from the other side: -6=2x I like to flip the equation so the x is on the left side, but that's just me: 2x=-6 Then divide the x: 2x/2=x and -6/2=3 Answer: x=3